OpenStudy (anonymous):

Would the ordered pair for 6m + n = 34 and m - 5n = 47 be (-8, 7) or (7, -8)?

OpenStudy (anonymous):

Multiply the first equation by 5 and add it to the second to get m, then sub back for n.

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

Let's go step by step. Multiply the first equation by 5.

OpenStudy (anonymous):

6(5) + n = 34 30 + n = 34

OpenStudy (anonymous):

This equation 6m + n = 34 Please multiply it by 5.

OpenStudy (anonymous):

I am having major connection problems

OpenStudy (anonymous):

That's Ok, I will wait.

OpenStudy (anonymous):

I do not understand how to multiply that equation by 5????

OpenStudy (anonymous):

Ummm..

OpenStudy (anonymous):

Yea, I know, it's pathetic...